# CompBench: Benchmarking Complex Instruction-guided Image Editing

### v2

@@ Abstract
While real-world applications increasingly demand intricate scene manipulation, existing instruction-guided image editing benchmarks often oversimplify task complexity and lack comprehensive, fine-grained instructions. To bridge this gap, we introduce CompBench, a large-scale benchmark specifically designed for complex instruction-guided image editing. CompBench features challenging editing scenarios that incorporate fine-grained instruction following, spatial and contextual reasoning, thereby enabling comprehensive evaluation of image editing models’ precise manipulation capabilities. To construct CompBench, We propose an MLLM-human collaborative framework with tailored task pipelines. Furthermore, we propose an instruction decoupling strategy that disentangles editing intents into four key dimensions: location, appearance, dynamics, and objects, ensuring closer alignment between instructions and complex editing requirements. Extensive evaluations reveal that CompBench exposes fundamental limitations of current image editing models and provides critical insights for the development of next-generation instruction-guided image editing systems. The dataset, code, and models are available in comp-bench.github.io.
Figure 1: Examples of CompBench. The figure showcases diverse instruction-guided image editing tasks across nine categories: object addition, object removal, object replacement, multi-object editing, multi-turn editing, implicit reasoning, action editing, location editing and viewpoint editing).
@@ 1 Introduction
Recent advances in instruction-guided image editing have pursued user-friendly and efficient manipulation of visual content. While such systems aim to simplify complex editing workflows, real-world applications often demand intricate instructions including spatial relationships, appearance details, and implicit reasoning. This necessitates the development of models with comprehensive capabilities in visual grounding, contextual understanding, and complex reasoning, thereby presenting substantial challenges to existing methodologies. However, as demonstrated in Figure 2, existing instruction-guided image editing benchmarks, e.g., Emu Edit [29], MagicBrush [38], and ReasonEdit [14], exhibit critical limitations in assessing these essential capabilities, primarily in three aspects:
Lack of Scene Complexity. A key limitation of existing benchmarks is their insufficient scene complexity, which hampers the representation of the intricate visual structures inherent in real-world images. This limitation primarily stems from two factors: the prevalent use of synthetic images and the selective inclusion of oversimplified real images.
The first issue arises from the widespread use of text-to-image generation models, such as Stable Diffusion [27], in previous benchmark construction [41, 22]. Although convenient, the resulting synthetic scenes typically exhibit sparse spatial layouts, limited foreground object diversity, minimal occlusions, and overly simplified textures and lighting conditions.
Such artificial compositions lack the dense object interactions, natural clutter, and photorealistic qualities essential for evaluating practical editing capabilities.
Even when incorporating real images from datasets, such as COCO [19], current benchmarks often present oversimplified scenarios. Although more realistic than synthetic counterparts, these images frequently exhibit elementary compositions that are insufficient for evaluating models on complex spatial relationships and interactions among multiple objects.
@@
This problem is further exacerbated by benchmark design choices, wherein creators often deliberately exclude highly complex scenes featuring heavy occlusions, intricate details, or dynamic elements due to the challenges they pose for ground truth construction. While this practice facilitates more controllable evaluation, it creates a concerning discrepancy between benchmark performance and real-world applicability. Consequently, image editing models may attain high metric scores on these relatively simplified benchmarks, yet remain inadequate for real-world editing tasks that demand advanced scene understanding and nuanced manipulation. For instance, in reasoning-based tasks, InstructPix2Pix [2] exhibits a notable performance decline on our CompBench compared with ReasonEdit [14], showing decreases of approximately 2.5 in PSNR, 0.02 in SSIM, and 0.4 in CLIP-Score.
@@
Limited Instruction and Task Comprehensiveness. Beyond their oversimplified visual scenes, current benchmarks are further constrained by the narrow scope of editing instructions and tasks, failing to reflect the complexity of real-world user demands. Most existing datasets rely on simplistic, atomic-level instructions (e.g., “change the dog to a cat”) that lack contextual reasoning, and compositional logic typical of real user requests. In practice, editing instructions often involve: multi-object manipulations (“remove the dog on the left and the cat behind the chair”), spatial relationships (“add a man to the right front of the woman”), or dynamic scene reasoning (“make the man in white bend down more”). The absence of such complex instructions artificially inflates model’s performance in perception tasks, as systems are not rigorously tested on the full spectrum of challenges that may be encountered in real-world applications. Furthermore, existing benchmarks present the deprivation in task diversity. They rarely cover multi-object editing scenarios, where multiple entities must be manipulated simultaneously under specific relational or contextual constraints (e.g., “remove the man in white and the woman behind the man”). Those benchmarks also tend to ignore action editing tasks, where models are required to adjust dynamic states or actions rather than static attributes (e.g., “make the dog jump higher”). Without assessing these realistic and complex editing scenarios, they provide an incomplete and often misleading evaluation of model robustness and practical applicability.
@@
Table 1: Comparison of existing image-editing datasets and benchmarks. Our benchmark supports seven core editing tasks, including multi-object editing, action editing and viewpoint editing, which are absent from most prior benchmarks. Scenario complexity is quantified by four indicators: Avg. Obj. (average number of objects per image), Avg. Cat. (average number of object categories per image), OCC (percentage of images that contain occluded objects), and OOF (percentage of images that contain out-of-frame objects). Across all four metrics, our benchmark exhibits the highest complexity, underscoring its suitability for rigorous evaluation.
@@
Datasets / Benchmarks Size Types Task Complexity Local Multi-turn Multi-obj. Implicit Action Location Viewpoint Avg. Obj. Avg. Cat. Occ. Rate OOF Rate Datasets InstructPix2Pix [2] 313K 4 8.71 4.16 79.36 81.39 EditWorld [37] 8.6K 1 8.01 4.45 76.67 72.00 UltraEdit [45] 4M 9 7.68 4.70 75.30 78.10 SEED-Data-Edit [10] 3.7M 6 6.21 3.82 63.82 81.40 HQ-Edit [15] 197K 6 8.22 4.84 66.97 60.30 AnyEdit [41] 2.5M 25 6.95 4.37 60.45 57.20 Benchmarks MagicBrush [38] 10K 5 9.22 5.04 91.71 78.34 EMU_Edit [29] 8 8.38 5.19 78.51 83.60 Reason-Edit [14] 0.2K - 4.93 3.09 54.30 52.28 I2EBench [22] 2K 16 7.03 4.20 68.78 66.40 Ours* 3K 9 13.58 5.87 98.47 86.38
@@
Figure 2: Comparison between current datasets or benchmarks and our CompBench. First row: failed cases of other benchmarks. These results fail to maintain background consistencies or introduce noticable artifacts into the editing region. Second row: Examples of other benchmarks. These cases lack scene complexity and instruction comprehensiveness. Third row: Examples of our CompBench. Our benchmark features complex real-world scenarios with precise instructions.
@@
Deficiencies in Edited Image Quality. Another critical yet often overlooked limitation of current benchmarks lies in the suboptimal quality of their edited images. Many existing datasets exhibit two predominant issues that compromise their reliability: (1) instruction-alignment inaccuracies, where the edited output fails to precisely fulfill the specified modifications. (2) conspicuous visual artifacts, including geometric distortions, contextual inconsistencies in backgrounds, and semantic incoherencies in foreground objects. These quality deficiencies introduce substantial noise into performance evaluations, potentially leading to misleading assessments of model capabilities. Such artifacts raises concerns about whether current benchmarks can effectively discriminate between truly sophisticated editing systems and those that merely produce superficially plausible results.
@@
To address the aforementioned issues, we introduce CompBench, the first large-scale benchmark for instruction-based image editing in complex scenarios, specific examples are illustrated in Figure 1. Our benchmark offers the following three major advantages:
@@
Realistic and Complex Scene Composition. As shown in Table 1, Our benchmark encompasses scenes that embody the diverse and realistic complexities present in real-world settings. We compare CompBench with existing datasets and benchmarks, including InstructPix2Pix [2], MagicBrush [38], Emu-Edit [29], Reason-Edit [14], Edit-World [37], I2E-Bench [22], UltraEdit [45], SEED-Data-Edit [10], HQ-Edit [15], and AnyEdit [41]–across four dimensions: average number of objects, average number of object categories, overall object occlusion rate, and out-of-frame object rate. CompBench consistently surpasses prior benchmarks in all these metrics. Notably, our average number of objects per image is approximately 47.3% higher than the second best (MagicBrush [38]), demonstrating the heightened complexity and diversity of our scenes.
@@
Comprehensive Task Coverage and High Difficulty Level. As depicted in Figure 4(a), CompBench encompasses five major categories, consisting of local editing, multi-editing, action editing, scene spatial editing, and complex reasoning, spanning a total of nine tasks. We provide detailed breakdowns of the six core capabilities these tasks require and analyze the degree to which our benchmark challenges each, which can be found in Appendix. Additionally, we propose an Instruction Decomposition Strategy to improve the clarity and precision of image editing instructions. Specifically, we introduce a structured framework that organizes editing instructions along four complementary dimensions: spatial positioning (e.g., “left of the table”), visual attributes (such as color or texture), motion states (e.g., “flying”), and object entities. This structured approach converts potentially ambiguous requirements into well-defined specifications without sacrificing the natural expressiveness of instructions. By systematically covering each aspect of an editing operation while preserving the flexibility of natural language, our framework produces instructions that are both intuitively understandable and technically precise for complex image editing tasks.
@@
High-Quality Data Curation. Every sample in CompBench is meticulously constructed through multiple rounds of expert review, ensuring the highest quality of edits. Unlike other benchmarks where editing failures are common, all data in CompBench represent successfully executed editing results, with SSIM (Structural Similarity Index Measure) scores significantly outperforming those of other datasets, as illustrated in Figure 4(b). This rigorous quality control ensures that CompBench provides a reliable assessment of model performance in realistically complex editing scenarios.
2 Related Work
@@
Instruction-guided Image Editing. Instruction-guided image editing enables efficient image manipulation using only textual editing instructions, without the need for manual mask drawing or providing explicit visual elements for editing. This approach aligns more closely with user intentions. Diffusion Models [11], particularly Stable Diffusion [27] (SD), facilitate this task significantly by supporting explicit text inputs. The advent of methods such as InstructPix2Pix [2], which are based on Diffusion Models [11], has greatly improved the effectiveness of instruction-based image editing. InstructPix2Pix [2] leverages large language models (LLMs) [32, 4, 3, 31] and text-to-image (T2I) [26, 25, 28, 27] models to generate large-scale datasets and trains a diffusion model that is capable of following natural language instructions. HIVE [44] introduces a reward model that leverages human feedback to align image editing outputs with human preferences. Approaches such as SmartEdit [14], MGIE [9], and the latest Step1X-Edit [21] interactively integrate image and instruction representations using multi-modal large language models (MLLMs) [18, 1, 20, 34], injecting these capabilities into diffusion models to better understand complex instructions and achieve more precise edits. AnyEdit [41] constructs an extremely large-scale multi-task dataset and adopts a mixture-of-experts (MoE) [8, 6] architecture to better accommodate diverse editing tasks. SEED-X [10] utilizes a visual tokenizer to unify image comprehension and generation, establishing a unified multi-granularity comprehension and generation model that enhances editing performance. GoT [7] employs Generation Chain-of-Thought [36] reasoning, incorporating stepwise reasoning into visual generation and editing processes, and allows users to explicitly follow the editing reasoning process, thereby enabling more refined edits.
@@
Image Editing Benchmarks.  High-quality image editing datasets and benchmarks are crucial for both model training and evaluation. MagicBrush [38] manually constructs a 10K-scale instruction-based image editing dataset, encompassing single-turn, multi-turn, mask-provided, and mask-free editing tasks. EMU-edit [29] introduces a challenging benchmark comprising seven diverse editing tasks. HQ-Edit [15] employs a scalable data collection pipeline to create a high-quality dataset of 200K instruction-guided image editing samples. SmartEdit [14] introduces Reason-Edit, a small-scale, manually curated benchmark focused on complex instruction-based image editing. Edit-world [37] presents the concept of world-instructed image editing and creates a dataset featuring instructions in a world context. I2EBench [22] proposes a comprehensive benchmark for instruction-based image editing evaluation, including automated multi-dimensional assessment. UltraEdit [45] develops a scalable framework for producing large and high-quality image editing datasets, introducing a large-scale instruction-based dataset. SEED-Data-Edit [10] provides a hybrid dataset composed of auto-generated, real-world, and human-annotated multi-turn editing samples, demonstrating its effectiveness in advancing instruction-based image editing.
@@ 3.1 Task Categorization and Definitions
Our complex instruction-guided image editing benchmark, CompBench, contains 3k+ image-instruction pairs. To enhance the comprehensiveness of evaluation, we categorize editing tasks into five major classes with nine specific tasks based on their characteristics:(1) Local Editing: focuses on manipulating local objects, including object removal, object addition and object replacement. (2) Multi-editing: addresses interactions among multiple objects or editing steps, including multi-turn editing and multi-object editing. (3) Action Editing: modifies the dynamic states or interactions of objects. (4) Scene Spatial Editing: alters scene spatial properties, consisting of location editing and viewpoint editing. (5) Complex Reasoning: requires implicit logical reasoning, including implicit reasoning. Examples of these tasks can be found in Figure 1.

### v4

@@ Abstract
While real-world applications increasingly demand intricate scene manipulation, existing instruction-guided image editing benchmarks often oversimplify task complexity and lack comprehensive, fine-grained instructions. To bridge this gap, we introduce CompBench, a large-scale benchmark specifically designed for complex instruction-guided image editing. CompBench features challenging editing scenarios that incorporate fine-grained instruction following, spatial and contextual reasoning, thereby enabling comprehensive evaluation of image editing models’ precise manipulation capabilities. To construct CompBench, we propose an MLLM-human collaborative framework with tailored task pipelines. Furthermore, we propose an instruction decoupling strategy that disentangles editing intents into four key dimensions: location, appearance, dynamics, and objects, ensuring closer alignment between instructions and complex editing requirements. Extensive evaluations reveal that CompBench exposes fundamental limitations of current image editing models and provides critical insights for the development of next-generation instruction-guided image editing systems. Our project page is available at https://comp-bench.github.io/.
@@ 1 Introduction
Figure 1: Examples of CompBench. The figure showcases nine tasks in our CompBench: object addition, object removal, object replacement, multi-object editing, multi-turn editing, implicit reasoning, action editing, location editing and viewpoint editing.
@@
Recent advances in instruction-guided image editing have pursued user-friendly and efficient manipulation of visual content. While such systems aim to simplify complex editing workflows, real-world applications often demand intricate instructions including spatial relationships, appearance details, and implicit reasoning. This necessitates the development of models with comprehensive capabilities in visual grounding, contextual understanding, and complex reasoning, thereby presenting substantial challenges to existing methodologies. However, as demonstrated in Figure 2, existing instruction-guided image editing benchmarks [30, 44, 14] exhibit critical limitations in assessing these essential capabilities, primarily in three aspects:
Lack of Scene Complexity.
A key limitation of current benchmarks is their insufficient scene complexity, which hampers the representation of intricate visual structures inherent in real-world images. Specifically, recent benchmark constructions [42, 22] predominantly source their images from general-purpose datasets, such as MS COCO [19]. While these benchmarks utilize real images, they often present oversimplified, object-centric scenarios with elementary compositions. These scenes typically feature sparse spatial layouts, limited foreground object diversity, and minimal occlusions, lacking the dense object interactions and natural clutter essential for evaluating practical editing capabilities. Consequently, they remain insufficient for comprehensively evaluating models on complex spatial relationships and interactions among multiple objects.
@@
This problem is further exacerbated by benchmark design choices, wherein creators often deliberately exclude highly complex scenes featuring heavy occlusions, intricate details, or dynamic elements due to the challenges they pose for ground truth construction. While this practice facilitates more controllable evaluation, it creates a concerning discrepancy between benchmark performance and real-world applicability.
Consequently, image editing models may attain high metric scores on these relatively simplified benchmarks, yet remain inadequate for real-world editing tasks that demand advanced scene understanding and manipulation. For instance, in reasoning-based tasks, InstructPix2pix [3] exhibits a notable performance decline on our CompBench compared with ReasonEdit [14], showing decreases of approximately 2.5 in PSNR, 0.02 in SSIM, and 0.4 in CLIP-Score.
Limited Instruction and Task Comprehensiveness.
@@
Beyond their oversimplified visual scenes, current benchmarks are further constrained by the narrow scope of editing instructions and tasks, failing to reflect the complexity of real-world user demands. Most existing datasets rely on simplistic, atomic-level instructions (e.g., “change the dog to a cat”) that lack contextual reasoning, and compositional logic typical of real user requests. In reality, user instructions often require complex reasoning and manipulation. These include multi-object editing (“remove the dog and the cat”), edits based on spatial relationships (“add a man to the right of the woman”), or action editing that modifies dynamic states (“make the man in white bend down more”). Current benchmarks, however, largely neglect these sophisticated task categories. This deficiency in instruction and task diversity prevents models from being rigorously tested on the full spectrum of challenges encountered in real-world applications. Consequently, their performance can be artificially inflated on simple tasks, providing an incomplete and misleading evaluation of true robustness and practical applicability.
@@
Table 1: Comparison of existing image-editing datasets and benchmarks. Our benchmark supports seven core editing tasks, including multi-object, action and viewpoint editing, which are absent from most prior benchmarks. Scenario complexity is quantified by four indicators: Avg. Obj. (average number of objects per image), Avg. Cat. (average number of object categories per image), OCC (percentage of images that contain occluded objects), and OOF (percentage of images that contain out-of-frame objects). Across all four metrics, our benchmark exhibits the highest complexity, underscoring its suitability for rigorous evaluation.
@@
Datasets / Benchmarks Size Types Task Complexity Local Multi-turn Multi-obj. Implicit Action Location Viewpoint Avg. Obj. Avg. Cat. Occ. Rate OOF. Rate Datasets InstructPix2pix [3] 313K 4 8.71 4.16 79.36 81.39 EditWorld [43] 8.6K 1 8.01 4.45 76.67 72.00 UltraEdit [48] 4M 9 7.68 4.70 75.30 78.10 SEED-Data-Edit [12] 3.7M 6 6.21 3.82 63.82 81.40 HQ-Edit [15] 197K 6 8.22 4.84 66.97 60.30 AnyEdit [42] 2.5M 25 6.95 4.37 60.45 57.20 ImgEdit [41] 1.2M 13 9.01 4.72 69.65 69.14 Benchmarks MagicBrush [44] 10K 5 9.22 5.04 91.71 78.30 EMU_Edit [30] 8 8.38 5.19 78.51 83.60 Reason-Edit [14] 0.2K - 4.93 3.09 54.30 52.28 I2EBench [22] 2K 16 7.03 4.20 68.78 66.40 GEdit-Bench [21] 0.6K 11 9.96 4.93 67.67 65.40 Complex-Edit [40] 1K 24 9.23 4.77 78.29 72.98 RefEdit [24] 20K 5 9.74 5.26 91.02 69.00 KRIS-Bench [39] 1.3K 22 6.04 3.09 29.49 40.69 ComplexBench-Edit [34] 763 10 7.85 4.75 75.71 74.14 Ours* 3K 9 13.58 5.87 98.47 86.38
@@
Figure 2: Comparison between current datasets or benchmarks and our CompBench. First row: failed cases of other benchmarks. These results fail to maintain background consistencies or introduce noticeable artifacts into the editing region. Second row: Examples of other benchmarks. These cases lack scene complexity and instruction comprehensiveness. Third row: Examples of our CompBench. Our benchmark features complex real-world scenarios with precise instructions.
Deficiencies in Edited Image Quality.
@@
Another critical limitation of current benchmarks is the suboptimal quality of their edited images. Many existing datasets exhibit two predominant issues that compromise their reliability: (1) instruction-alignment inaccuracies, where the edited output fails to precisely fulfill the specified modifications. (2) conspicuous visual artifacts, such as geometric distortions, background inconsistencies, or semantically incoherent objects. These quality deficiencies introduce substantial noise into performance evaluations, potentially leading to misleading assessments of model capabilities. Consequently, such benchmarks may fail to effectively discriminate between truly sophisticated editing systems and those that merely produce superficially plausible but flawed results.
@@
To address the aforementioned issues, we introduce CompBench, the first large-scale benchmark for instruction-guided image editing in complex scenarios, specific examples are illustrated in Figure 1. Our benchmark offers the following three major advantages:
Realistic and Complex Scene Composition.
@@
As shown in Table 1, our benchmark encompasses scenes that embody the diverse and realistic complexities present in real-world settings. We compare CompBench with existing datasets and benchmarks across four dimensions: average number of objects, average number of object categories, overall object occlusion rate, and out-of-frame object rate. Details of these metrics are shown in the supplementary material. CompBench consistently surpasses prior benchmarks in all these metrics. Notably, our average number of objects per image is approximately 36.3% higher than the second best (GEdit-Bench [21]), demonstrating the heightened complexity and diversity of our scenes.
Comprehensive Task Coverage and High Difficulty Level.
@@
As depicted in Figure 4, CompBench encompasses five major categories, consisting of local editing, multi-editing, action editing, scene spatial editing, and complex reasoning, spanning a total of nine tasks. These tasks are designed to challenge six core capabilities, with a detailed analysis of our benchmark’s difficulty for each provided in the supplementary material. Additionally, we propose an Instruction Decomposition Strategy to improve the clarity and precision of image editing instructions. Specifically, we structure editing instructions along four dimensions: spatial positioning (e.g., “left of the table”), visual attributes (such as color or texture), motion states (e.g., “flying”), and object entities. This structured approach converts potentially ambiguous requests into well-defined specifications without sacrificing the natural expressiveness of instructions. By systematically covering each aspect of an editing operation while preserving the flexibility of natural language, our method produces instructions that are both intuitively understandable and technically precise for complex image editing tasks.
High-Quality Data Curation.
@@
Every sample in CompBench is meticulously constructed through multiple rounds of expert review, ensuring the highest quality of edits. Unlike other benchmarks where editing failures are common, all data in CompBench represent successfully executed editing results, with SSIM (Structural Similarity Index Measure) scores significantly outperforming those of other datasets, as illustrated in Figure 5. This rigorous quality control ensures that CompBench provides a reliable assessment of model performance in realistically complex editing scenarios.
2 Related Works
Instruction-guided Image Editing.
@@ 2 Related Works
Instruction-guided image editing enables efficient image manipulation using only textual editing instructions, eliminating the need for manual mask or explicit visual inputs and better aligning with user intent. Diffusion models [13], particularly Stable Diffusion [28] (SD), facilitate this task significantly by supporting explicit text inputs. Methods built upon diffusion models such as InstructPix2pix [3], have greatly improved editing effectiveness. InstructPix2pix leverages large language models (LLMs) [33, 6, 4, 32] and text-to-image (T2I) [26, 27, 28, 29] models to generate large-scale datasets and trains a diffusion model that is capable of following natural language instructions. HIVE [47] introduces a reward model that leverages human feedback to align edits with human preferences. Approaches such as SmartEdit [14], MGIE [11], and Step1X-Edit [21] integrate image and instruction representations using multi-modal large language models (MLLMs) [18, 1, 20, 35], injecting these capabilities into diffusion models for more precise control. AnyEdit [42] constructs an extremely large-scale multi-task dataset and adopts a mixture-of-experts (MoE) [10, 8] architecture to better accommodate diverse editing tasks. SEED-X [12] utilizes a visual tokenizer to unify image comprehension and generation, establishing a unified multi-granularity comprehension and generation model that enhances editing performance. GoT [9] incorporates Generation Chain-of-Thought [37] reasoning into the editing process, allowing for more refined, step-by-step edits. Recently, FLUX.1 Kontext [16] applies flow matching to build a unified image generation and editing model. Bagel [5] adopts a decoder only architecture to construct a multimodal understanding and generation model. Qwen-Image-Edit [38], the editing model of Qwen-Image [38], demonstrates strong text rendering and image editing capabilities.
Image Editing Benchmarks.
@@
High-quality image editing datasets and benchmarks are crucial for model training and evaluation. Several notable benchmarks have been proposed: MagicBrush [44] provides a manually curated 10K dataset covering single-turn, multi-turn, mask-provided, and mask-free editing tasks. EMU-edit [30] introduces a challenging benchmark comprising seven diverse editing tasks. HQ-Edit [15] employs a scalable data collection pipeline to create a high-quality dataset of 200K instruction-guided image editing samples. SmartEdit [14] introduces Reason-Edit, a small-scale, manually curated benchmark focused on complex instruction-based image editing. Edit-world [43] presents the concept of world-instructed image editing and creates a dataset featuring instructions in a world context. I2EBench [22] proposes a comprehensive evaluation benchmark with automated multi-dimensional assessment. UltraEdit [48] develops a scalable framework for producing large and high-quality image editing datasets, introducing a large-scale instruction-based dataset. SEED-Data-Edit [12] provides a hybrid dataset composed of auto-generated, real-world, and human-annotated multi-turn editing samples. More recently, ImgEdit [41] introduces a large scale image editing dataset and a benchmark with multiple aspects. Step1X-Edit [21] construct GEdit-Bench [21] featuring real-world user instructions. Complex-Edit [40] adopts a “Chain-of-Edit” pipeline to develop an image editing benchmark across instructions of different complexity. ComplexBench-Edit [34] addresses combinatorial reasoning challenges by introducing a benchmark for chain-dependent instructions and a region-specific consistency metric. KRIS-Bench [39] evaluates the cognitive capabilities of models through a knowledge-based taxonomy spanning factual, conceptual, and procedural dimensions. Furthermore, RefEdit [24] targets the precise editing of specific objects in complex, multi-entity scenes based on referring expressions.
@@ 3.1 Task Categorization and Definitions
Our complex instruction-guided image editing benchmark, CompBench, contains over 3k image-instruction pairs. For comprehensive evaluation, we categorize the tasks into five major classes encompassing nine specific tasks: (1) Local Editing: manipulating objects via removal, addition, or replacement. (2) Multi-editing: addressing interactions across multi-turn or multi-object editing. (3) Action Editing: modifying dynamic states or object interactions. (4) Scene Spatial Editing: altering spatial properties through location or viewpoint editing. (5) Complex Reasoning: performing implicit contextual edits that require logical reasoning. Examples are illustrated in Figure 1.